OpenStudy (e_s_j_f):

hey, can someone please help me with with question. i need to integrate p = density, v = velocity, -k is constant.

OpenStudy (e_s_j_f):

\[\int\limits_{}^{} p v \frac{ dv }{ dx } = - k\]

OpenStudy (e_s_j_f):

@ganeshie8

OpenStudy (e_s_j_f):

@dan815

OpenStudy (anonymous):

can i take a total guess?

OpenStudy (e_s_j_f):

okay..

OpenStudy (anonymous):

will you know if its correct or not?

OpenStudy (e_s_j_f):

that's the thing. i've done it but i don't know if its correct.

OpenStudy (e_s_j_f):

\[v(x) = -kpvx\] seems wrong to me

OpenStudy (e_s_j_f):

\[v(x) = -kp \frac{ v^2 }{ 2 }x\]

OpenStudy (dan815):